OpenStudy (thedj4jc):

3x + 5y = 78 2x - y = 0 The point of intersection of the lines has an x-coordinate of _____. 78 6 -6 Please explain intersections to me, cause they make about zero sense.

OpenStudy (solomonzelman):

you can multiply the 2nd equation times 5 and add the equations.

OpenStudy (solomonzelman):

or, y=2x and sub

OpenStudy (texaschic101):

I was gonna say that

OpenStudy (texaschic101):

lol

OpenStudy (solomonzelman):

I figured. I was kind of "proposing more complicated approach" (complicated, comparing to y=2x sub)

OpenStudy (solomonzelman):

Dj, which was do you want to go ? Substitution or elimination ?

OpenStudy (thedj4jc):

The elimination sense to make more sense to me usually, but either works.

OpenStudy (solomonzelman):

yes, so we will do substitution if elimination is something you comprehend. (To get all concepts going)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 3x + 5y = 78 \\[0.4em] }\) (1st equation) \(\large\color{black}{ \displaystyle 2x - y = 0 }\) (2nd equation)

OpenStudy (solomonzelman):

Add y to both sides in the 2ND equation. What will your new (second) equation now be?

OpenStudy (thedj4jc):

2x = 0+y, right?

OpenStudy (solomonzelman):

yes, so that gives a second equation of y=2x (right ? )

OpenStudy (solomonzelman):

OUR NEW SYSTEM NOW IS \(\large\color{black}{ \displaystyle 3x + 5y = 78 \\[0.4em] }\) (1st equation) \(\large\color{black}{ \displaystyle y=2x }\) (2nd equation)

OpenStudy (thedj4jc):

Ok

OpenStudy (solomonzelman):

Now, what does the second equation really say? It tells you that [in this system of equations] "y" is the same thing (or, 'has the same value as') as "2x". that means that you can SUBSTITUTE "2x" instead of "y", into the 1st equation.

OpenStudy (solomonzelman):

Our first equation. \(\large\color{black}{ \displaystyle 3x + 5\color{red}{(y)} = 78 \\[0.4em] }\) Substitution \(\large\color{black}{ \displaystyle 3x + 5\color{red}{(2x)} = 78 \\[0.4em] }\)

OpenStudy (solomonzelman):

is this substitution making sense ?

OpenStudy (thedj4jc):

Not yet, but it's starting to I think.

OpenStudy (solomonzelman):

sure, if you read of the replies and have any question, make sure to ask.

OpenStudy (solomonzelman):

read **over** the replies.... (typo correction)

OpenStudy (thedj4jc):

I'm still not getting it. It makes sense that 2x = y, but I don't get how to finish or understand it.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y }\) is same thing as \(\large\color{black}{ \displaystyle 2x }\) (according to our 2nd equation) So the first equation that says \(\large\color{black}{ \displaystyle 3x+5\color{red}{y}=78 }\) would be same exact thing as \(\large\color{black}{ \displaystyle 3x+5(\color{red}{2x})=78 }\)

OpenStudy (solomonzelman):

by doing this substitution we are not changing the value of the equation, just re-writing in the second term (the "5y") as "5 times 2x". This allows us to find the X, but having to solve a simple equation with one variable.

OpenStudy (thedj4jc):

Ok, that makes sense. So, how do you find x now?

OpenStudy (solomonzelman):

well, you expand the \(\large\color{black}{ \displaystyle 5(\color{red}{2x}) }\) part, add all x's and on...

OpenStudy (solomonzelman):

5(2x) => 10x correct ?

OpenStudy (thedj4jc):

Yeah, so that would leave 3x + 10x, which could be simplified to 13x?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

So you get 13x=78

OpenStudy (solomonzelman):

and then x = ?

OpenStudy (thedj4jc):

x = 6?

OpenStudy (solomonzelman):

yup

OpenStudy (thedj4jc):

Thank you so much!!! This makes so much sense now!

OpenStudy (solomonzelman):

very nice. we have found that x=6.... remember we said that y=2x ?

OpenStudy (thedj4jc):

So, y = 12

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

any questions ?

OpenStudy (thedj4jc):

Nope, I think I've got it!! Thank you!!

OpenStudy (solomonzelman):

You welcome !